# Training-free LLM Verification via Recycling Few-shot Examples

### v1

@@ 4.2 Main results
Table 1: Main Results. Overall performance on seven reasoning benchmarks comparing the proposed ReFeri with different baselines not require additional training, under three different state-of-the-art LLMs. The best and second-best scores are highlighted in bold and underline, respectively.
@@
Models Methods MATH500 MMLU-pro GPQA DROP HotpotQA MuSR-op MuSR-ta Avg.
(Acc.) (Acc.) (Acc.) (EM / F1) (EM / F1) (Acc.) (Acc.)
GPT-4o-mini Zero-shot CoT 76.4 63.0 43.0 77.6 / 85.6 31.5 / 41.4 58.1 56.2 58.0
Few-shot CoT 75.2 63.0 41.3 76.8 / 83.1 34.0 / 45.1 59.4 77.0 61.0
USC 78.6 62.5 42.4 78.8 / 85.8 36.6 / 48.2 59.8 76.4 62.2
CoT-WP 77.8 64.2 42.4 77.6 / 83.1 34.0 / 45.6 57.0 79.6 61.8
LEAP 74.5 63.2 43.9 75.8 / 83.0 34.0 / 45.1 59.8 74.4 60.8
ReFeri (Ours) 78.2 65.0 42.4 79.6 / 85.3 36.2 / 47.9 61.3 82.8 63.6
GPT-4o Zero-shot CoT 77.5 73.9 48.8 75.1 / 85.3 37.6 / 49.9 61.7 66.6 63.0
Few-shot CoT 75.6 73.7 47.8 80.6 / 89.2 44.6 / 58.4 69.7 87.0 68.4
USC 79.8 72.1 50.5 82.0 / 90.2 45.8 / 60.4 73.4 87.6 70.2
CoT-WP 78.0 75.0 48.5 83.0 / 90.8 48.0 / 61.2 70.3 88.8 70.2
LEAP 75.6 74.0 45.5 81.5 / 89.8 45.1 / 58.4 66.8 87.2 68.0
ReFeri (Ours) 78.4 75.5 51.5 83.6 / 91.2 47.4 / 61.0 72.3 90.8 71.4
LLaMA-3.1-8B Zero-shot CoT 44.2 39.8 21.6 60.4 / 66.4 15.2 / 21.2 50.6 43.0 39.3
Few-shot CoT 42.9 38.7 24.0 61.4 / 67.3 19.0 / 25.1 53.3 64.8 43.1
USC 49.6 35.6 28.8 69.6 / 75.8 24.4 / 32.5 52.3 67.2 46.8
CoT-WP 47.8 44.8 32.3 70.2 / 75.1 25.0 / 32.2 56.6 71.6 49.8
LEAP 42.3 37.3 27.8 58.2 / 64.1 19.9 / 26.8 51.6 69.2 43.8
ReFeri (Ours) 51.0 45.1 34.8 70.2 / 76.7 25.0 / 33.0 56.6 80.0 51.8
@@
Table 1 summarizes the experimental results across seven different reasoning benchmarks and three different LLMs. For instance, across all tested LLMs and benchmarks, ReFeri improves average accuracy by 4.8% over Few-shot CoT, which corresponds to apply random selection instead. Compared to the second-best method, CoT-WP, ReFeri achieves an average improvement of 1.7% across all benchmarks. Notably, CoT-WP relies solely on the forward likelihood of each candidate, while ReFeri combines both forward and backward signals via a Bayes-deriven scoring function. This bidirectional formulation allows ReFeri to capture not just the confidence of an answer, but also its consistency with few-shot examples upon the LLM’s intrinsic knowledge about the task; consequently, it enables a better selection across various tasks. We note that performance of prompt-base selection, USC, largely varies depending on the task and used LLMs, which reveals the limitation of solely relying on LLM’s intrinsic knowledge. In addition, as mentioned in Section 4.1, MuSR is a benchmark where few-shot examples play a critical role [27] and our results also support this with 21.0% average improvement by Few-shot CoT over Zero-shot CoT. Here, we find that ReFeri further enlarges the improvement with the largest gain, outperforming the second-best method by 4.5%. This result shows that ReFeri is particularly effective in new domains where there LLM has a little prior knowledge and need to heavily rely on a few examples without additional training or reward models.
@@
Next, to assess whether ReFeri scales effectively with the number of candidate outputs similar to the conventional reward-based best-of-$K$ selection, we evaluate performance as the candidate pool grows. Specifically, we test $K=\{1,5,10,15\}$ candidates on three representative tasks—MATH500, GPQA, and MuSR-ta by using GPT-4o-mini as the generation model under Few-shot CoT. Across the three tasks, ReFeri yields consistent improvements as $K$ increases. On MATH500, while the accuracy of random selection decreases as the number of generated samples increases, ReFeri consistently selects higher-quality responses, improving from 75.8% at $K=1$ to 79.2% at $K=15$. On GPQA, where ReFeri raises performance from 41.4% to 45.5% as the candidate pool grows. Consistently, the largest gain is observed on MuSR-ta, which saw a sharp jump in accuracy from 75.6% to 86.0%, an improvement of 10.4%. In contrast, while USC is effective on MATH500, its performance even degrades with more candidates on GPQA and MuSR-ta, where candidate responses are more diverse and LLM has relatively less relavant knowledge. In addition, we observe an inherent ordering bias in USC: over 90% of selections come from the first two responses (see Appendix B.3), indicating a limitation of this prompt-base approach. Overall, these results confirm that ReFeri scales well with more candidates, demonstrating effectiveness and reliability in practical test-time scaling.
@@
Figure 3: Test-time scaling with ReFeri. Accuracy of ReFeri versus other training-free selection methods (Random selection and USC) on MATH500, GPQA, and MuSR-ta. GPT-4o-mini generate different numbers of candidate responses ($K=1,5,10,15$) using Few-shot CoT, similar to Table 1. Solid lines indicate the performance of ReFeri (ours), USC, and Random, respectively.
Table 2: Ablation study. Evaluation of scoring variants on responses by GPT-4o-mini, comparing the contribution of metric term (forward, backward, and direct) on MATH500 and GPQA.
Forw. Back. Dire. MATH500 GPQA
75.2 41.3
77.6 41.9
75.2 41.9
77.8 42.4
ReFeri 78.2 42.4
Table 3: Different few-shot examples. Accuracy across three different choices of few-shot examples on MATH500 (top) and GPQA (bottom) using GPT-4o-mini to generate responses.
@@
Few-shot CoT 74.4 73.8 74.7 74.3
ReFeri (Ours) 76.8 77.4 79.0 77.7
Few-shot CoT 41.7 39.7 44.8 42.1
ReFeri (Ours) 45.5 41.4 46.0 44.3
@@ 4.3 Additional analyses
In this section, we conduct additional analyses of ReFeri. We mainly conduct experiments using GPT-4o-mini on MATH500, GPQA, and MuSR-ta. More results are presented in Appendix B.
@@
Ablation study. To better understand which components contribute to the effectiveness of ReFeri, we conduct an ablation study on each part of the proposed scoring function (Eq. 10). We focus on two primary components: the forward confidence score (Eq. 6) and the backward consistency score (Eq. 8), both of which are grounded in Bayes’ rule (Eq. 5). The results are presented in Table 2. First, it is observed that combining forward and backward scores outperforms either score alone, validating the joint consideration. This complementary effect is from their different nature; while forward score focuses on model-generated response which may contain noise, backward score utilizes given few-shot examples which are well-curated inputs and ground-truth labels, and thus relatively clean. Accordingly, the forward and backward scores anchored on few-shot examples form the structural backbone of our method. Interestingly, adding the direct estimation like Eq. 10 consistently improves accuracy across benchmarks, though it is not required by the Bayes formulation. This result suggests its utility as a supplementary prior for verifying LLM outputs. More results are in Appendix B.
@@
Estimation models and few-shot examples. To examine whether ReFeri is sensitive to the choice of estimation model $p_{\theta}$, we evaluate its performance using three LLMs with diverse scales and architectures: LLaMA-3.2-1B-Instruct [54], Qwen-2.5-7B-Instruct [55], and LLaMA-3.1-70B-Instruct [56]. The generation model is fixed (either GPT-4o-mini, GPT-4o, or LLaMA-3.1-8B), and we apply each estimation models to two tasks on MATH500 and GPQA. The average accuracy of three generation LLMs is presented in Figure 4 (Full results are in Appendix B). Here, ReFeri consistently improves Few-shot CoT across all settings, with an average gain of 5.0% on MATH500 and 5.3% on GPQA. Notably, the smallest model (LLaMA-3.2-1B) performs competitively, and even achieves competitive performance on MATH500. We attribute this to the relative simplicity of MATH benchmark, as recent small LLMs often exhibit reasonable performance; hence, they can make reliable likelihood estimates for selection. In contrast, GPQA requires more complex reasoning; therefore, using the large estimation model could be more beneficial. Indeed, LLaMA-3.1-70B achieves the best performance on this case. Despite these task-specific differences, the overall improvements are consistent across all estimation models. This suggests that the effectiveness of ReFeri primarily stems from its validation strategy with few-shot examples, rather than the specific choice of estimation model.
Also, ReFeri highly relies on few-shot examples for scoring of both forward and backward scores (Section 3.2). This raises the question of how sensitive the method is to the choice of few-shot exemplars. To this end, we conduct a sensitivity study on MATH500 and GPQA using GPT-4o-mini, where we randomly sample three different few-shot examples from the training dataset. As shown in Table 3, both Few-shot CoT and ReFeri show some variation across seeds. Nevertheless, ReFeri consistently outperforms Few-shot CoT which corresponeds to random selection, and the average gap remains approximately 2.8%. These results indicate that ReFeri remains robust to exemplar choice and is consistently effective, rather than overfitted to specific demonstrations.
Table 4: Generalization to reasoning LLM. Accuracy of DeepSeek-R1-Distill-Qwen-7B [8] under zero-shot CoT and few-shot CoT methods with or without ReFeri.
Methods AIME24 AMC23
Zero-shot CoT 52.7 84.5
+ ReFeri (Ours) 63.3 90.0
Few-shot CoT 44.0 78.0
+ ReFeri (Ours) 50.0 85.0
Generalization to reasoning LLM. As discussed in Section 2, few-shot prompting is often less effective for RL-tuned LLMs such as DeepSeek-R1 [8]. This degradation may stem from several factors, including sensitivity to prompt formatting, incompatibility with few-shot templates, or reliance on special tokens (e.g., <think>). We also observe similar results through experiments on two challenging benchmarks, AIME24 [57] and AMC23 [58], with DeepSeek-R1-Distill-Qwen-7B [59]; for the experiments, we generate five reasoning paths under a temperature of 1.0 with five few-shot examples from MATH500, following [47]. We use this model for both generation and estimation. More detailed setups are presented in Appendix A.3. As shown in Table 4, Zero-shot CoT achieves 52.7% and 84.5% respectively, while Few-shot CoT yields a notable drop to 44.0% (-8.7%) and 78.0% (-6.5%). Nevertheless, the proposed ReFeri successfully selects the promising reasoning path and yields the significant accuracy gain (+6.0% and +7.0%, respectively). This result confirms the robustness of ReFeri from the utilization of well formatted few-shot examples for the validation.
On the other hand, one can observe that the final accuracy is still lower than the average accuracy by Zero-shot CoT, mainly due to the limited accuracy of reasoning paths from Few-shot CoT. However, as described in Eq. 3, ReFeri is indeed applicable to select reasoning paths of zero-shot CoT, while we primarily apply to few-shot CoT as it usually yields better reasoning paths (Table 1). With the experiments, we verify that applying ReFeri to Zero-shot CoT yields substantial improvements that raise accuracy to 63.3% (+10.6%) on AIME24 and 90.0% (+5.5%) on AMC23, respectively. These results are evidence that few-shot examples in ReFeri mainly serve as a form of post-hoc validation pipeline, not the generation guidance like conventional Few-shot CoT. Also, this effectiveness of ReFeri under decoupling between generation and selection suggests a robust alternative to conventional few-shot prompting strategies, particularly in settings where few-shot examples are ineffective with LLMs.
Table 5: LLM personalization. Evaluation results on LaMP-4 and LaMP-5 using GPT-4o-mini as generator. Vanilla uses no history, while Few-shot RAG retrieves user history via BM25.
Methods LaMP-4 LaMP-5
Rouge-1 Rouge-L Rouge-1 Rouge-L
Vanilla 0.120 0.106 0.421 0.332
Few-shot RAG 0.142 0.126 0.453 0.368
ReFeri (Ours) 0.160 0.141 0.503 0.402
Application to LLM personalization. Lastly, we further apply ReFeri for LLM personalization to evaluate its broader applicability. The goal of LLM personalization is steering LLMs’ responses towards the individual users, which becomes progressively important [15, 60, 61]. One representative baseline for LLM personalization is few-shot retrieval-augmented generation (RAG) that retrieved the user’s previous data relevant to the given test query, and hence it’s natural to apply ReFeri. Specifically, we evaluate on two tasks in LaMP benchmark [15], LaMP-4 (personalized news headline generation) and LaMP-5 (personalized scholarly title generation), and use GPT-4o-mini as generation LLM. We generate $K=5$ candidate responses with a temperature of 1.0 as same as Table 1. Vanilla baseline directly answers to query without external context, while the Few-shot RAG baseline augments input prompt with $N=5$ examples retrieved via BM25 [62] from the user’s history. Following [15], we evaluate all responses against gold references using ROUGE-1 and ROUGE-L. The average of all $K$ responses is reported for the baselines, and results with the selected response is reported for ReFeri, respectively. As shown in Table 5, ReFeri consistently outperforms both baselines across LaMP-4 and LaMP-5. Notably, it improves ROUGE-L from 0.368 to 0.402 on LaMP-5, and from 0.126 to 0.141 on LaMP-4. This result demonstrates the applicability of ReFeri beyond traditional reasoning tasks—to open-ended, user-specific scenarios.

### v2

@@ 3.2 Main results
Table 1: Main results. Overall performance on seven reasoning benchmarks comparing the proposed ReFeri with different baselines not require additional training, under three different state-of-the-art LLMs. The best and second‑best scores are highlighted in bold and underline, respectively.
@@
Models Methods MuSR-ta MuSR-op GPQA MATH500 DROP HotpotQA MMLU-PRO Avg.
(Acc.) (Acc.) (Acc.) (Acc.) (EM / F1) (EM / F1) (Acc.)
LLaMA-3.1-8B
Zero-shot CoT 43.0 50.6 21.6 44.2 60.4 / 66.4 15.2 / 21.2 39.8 39.3
Few-shot CoT 64.8 53.3 24.0 42.9 61.4 / 67.3 19.0 / 25.1 38.7 43.4
LEAP 69.2 51.6 27.8 42.3 58.2 / 64.1 19.9 / 26.8 37.3 43.8
USC 67.2 52.3 28.8 49.6 69.6 / 75.8 24.4 / 32.5 45.6 48.2
CoT-WP 72.4 54.7 29.3 47.8 71.6 / 75.8 25.8 / 33.4 46.0 49.7
ReFeri 79.6 57.8 35.4 51.2 69.4 / 75.7 25.0 / 33.2 45.1 51.9
GPT-4o-mini
Zero-shot CoT 56.2 58.1 43.0 76.4 77.6 / 85.6 31.5 / 41.4 63.0 58.0
Few-shot CoT 77.0 59.4 41.3 75.2 76.8 / 83.1 34.0 / 45.1 63.0 61.0
LEAP 74.4 59.8 43.9 74.5 75.8 / 83.0 34.0 / 45.1 63.2 60.8
USC 74.4 60.9 46.0 77.8 76.8 / 83.8 35.0 / 47.2 63.7 62.1
CoT-WP 78.8 56.3 42.4 77.8 76.4 / 82.5 35.8 / 46.7 64.6 61.7
ReFeri 82.8 61.3 41.9 77.8 79.2 / 84.9 36.2 / 48.0 64.9 63.4
GPT-4o
Zero-shot CoT 66.6 61.7 48.8 77.5 75.1 / 85.3 37.6 / 49.9 73.9 63.0
Few-shot CoT 87.0 69.7 47.8 75.6 80.6 / 89.2 44.6 / 58.4 73.7 68.4
LEAP 87.2 66.8 45.5 75.6 81.5 / 89.8 45.1 / 58.4 74.0 68.0
USC 85.2 71.1 47.0 77.4 82.2 / 90.2 45.6 / 59.7 74.5 69.0
CoT-WP 88.0 68.8 47.5 78.4 83.4 / 91.4 47.2 / 60.2 74.1 69.6
ReFeri 90.4 71.9 51.5 77.8 83.6 / 91.1 47.0 / 60.7 75.4 71.1
@@
Table 1 summarizes the experimental results across seven different reasoning benchmarks and three different LLMs. For instance, across all LLMs and benchmarks, ReFeri improves average accuracy by 4.5% over Few-shot CoT, which corresponds to apply random selection instead. Compared to the second-best method, CoT-WP, ReFeri achieves an average improvement of 1.8% across all benchmarks. Notably, CoT-WP relies solely on the forward likelihood of each candidate, while ReFeri combines both forward and backward signals via a Bayes-derived scoring function. This bidirectional formulation allows ReFeri to capture not just the confidence of an answer, but also its consistency with few-shot examples upon the LLM’s intrinsic knowledge about the task; consequently, it enables a better selection across various tasks. We note that performance of prompt-based selection, USC, largely varies depending on the task and used LLMs, which reveals the limitation of solely relying on LLM’s intrinsic knowledge. In addition, as mentioned in Section 3.1, MuSR is a benchmark where few-shot examples play a critical role (Sprague et al., 2025) and our results also support this with 21.0% average improvement by Few-shot CoT over Zero-shot CoT. Here, we find that ReFeri further enlarges the improvement with the largest gain, outperforming the second-best method by 4.5%. This result shows that ReFeri is particularly effective in new domains where LLM has little prior knowledge and need to heavily rely on a few examples without additional training or reward models.
@@
Next, to assess whether ReFeri scales effectively with the number of candidate outputs similar to the conventional reward-based best-of-$K$ selection, we evaluate performance as the candidate pool grows. Specifically, we test $K=\{1,5,10,15,20\}$ candidates on three representative tasks—MATH500, GPQA, and MuSR-ta by using GPT-4o-mini as the generation model under Few-shot CoT. Across the three tasks, ReFeri yields consistent improvements as $K$ increases. On MATH500, while the accuracy of random selection decreases as the number of generated samples increases, ReFeri consistently selects higher-quality responses, improving from 75.8% at $K=1$ to 79.4% at $K=20$. On GPQA, where ReFeri raises performance from 41.4% to 45.5% as the candidate pool grows. Consistently, the largest gain is observed on MuSR-ta, which saw a sharp jump in accuracy from 75.6% to 86.0%, an improvement of 10.4%. In contrast, CoT-WP and USC exhibit unstable accuracy under the test-time scaling. Their performance even degrades as the number of candidates increases, suggesting that these methods do not capture what is truly plausible among the candidates. Notably, USC demonstrates strong performance on GPQA when K=5, but its accuracy declines as K increases, highlighting sensitivity to the candidate set size. In addition, we observe an inherent ordering bias in USC: selections come from the first two responses regardless of correctness (see Appendix B.1), indicating a limitation of prompt-based approach. Overall, these results confirm that ReFeri scales well with more candidates, demonstrating effectiveness and reliability in practical test-time scaling.
@@
Figure 3: Test-time scaling with ReFeri. Accuracy of ReFeri versus other training-free selection methods (Random selection, CoT-WP and USC) on MATH500, GPQA, and MuSR-ta. GPT-4o-mini generate different numbers of candidate responses ($K=1,5,10,15,20$) using Few-shot CoT.
Table 2: Ablation study. Evaluation of scoring variants averaged over three generation LLMs, comparing the contribution of metric term (forward and backward) on MATH500 and GPQA.
Forw. Back. MATH500 GPQA
64.6 37.7
68.9 42.2
64.6 34.7
ReFeri (Full) 69.0 42.7
ReFeri 68.9 42.9
Table 3: Different few-shot examples. Accuracy across three different choices of few-shot examples on MATH500 (top) and GPQA (bottotm) using GPT-4o-mini to generate responses.
@@
Few-shot CoT 75.2 74.5 75.0 74.9
ReFeri (Ours) 77.8 79.0 77.8 78.2
Few-shot CoT 41.3 41.5 38.9 40.6
ReFeri (Ours) 41.9 43.4 41.9 42.4
@@ 3.3 Additional analyses
In this section, we present additional analyses of ReFeri. We primarily perform experiments with GPT-4o-mini on MATH500 and GPQA datasets. More results are presented in Appendix B.
@@
Ablation study. To better understand which components contribute to the effectiveness of ReFeri, we conduct an ablation study on each part of the proposed scoring function (Eq. 11), which is grounded in Bayes’ rule (Eq. 5). In Table 2, we report performance on MATH500 and GPQA, averaged across the three LLMs used in Table 1. First, one can observe that using the combined score yields better results compared to solely using the forward confidence score (Eq. 6) or the backward consistency score (Eq. 10). This complementary effect comes from their different natures; while the forward score focuses on model-generated response which may contain noise, backward score utilizes given few-shot examples, which are well-curated inputs and ground-truth labels, and thus relatively clean. Next, it is also observed that using the cost-efficient variant of backward term (ReFeri) does not compromise the performance compared to ReFeri (Full) which uses original backward score (Eq. 8), when combined with forward term (more results are in Table 16). This result mitigates concerns regarding the additional computations incurred by original backward term.
Robustness to few-shot examples. ReFeri highly relies on few-shot examples for scoring of both forward and backward scores (Section 2.2). This raises the question of how sensitive the method is to the choice of few-shot examples. To answer this, we conduct a sensitivity study on MATH500 and GPQA using GPT-4o-mini, where we use three different few-shot examples with one original and two newly sampled. As shown in Table 3, both Few-shot CoT and ReFeri show variation across these different sets. Nevertheless, ReFeri consistently outperforms Few-shot CoT which corresponds to random selection, and the average gap remains 2.6%. These results indicate that ReFeri remains robust to exemplar choice and is consistently effective, rather than overfitted to specific demonstrations.
Moreover, in practical applications, the clarity of few-shot example might be not always guaranteed. To verify the effectiveness of ReFeri under this scenario, we first synthesize low-quality few-shot examples by converting the original examples via prompting GPT-4o-mini to degrade the quality of reasoning in data. The degradation of quality is indeed confirmed through LLM-as-judge framework (results and judgments are in Appendix B.4 ) As shown in Table 11, ReFeri maintains consistent improvements even under degraded exemplars, indicating that recycled few-shot examples as verification remains effective without well curated examples; for instance, on MATH500 with LLaMA-3.1-8B, accuracy improves from 39.5% to 47.6% (+8.1), demonstrating ReFeri’s robustness.
Table 4: Ablation on generation/evaluation prompts. Evaluation on MATH500 and GPQA with generation/evaluation prompt variants (orig/plan/role).
Gen Eval MATH500 GPQA
Few-shot ReFeri Few-shot ReFeri
Orig Orig 75.2 77.8 41.3 41.9
Plan 75.2 78.0 41.3 42.4
Role 75.2 77.8 41.3 41.9
Plan Plan 74.6 78.2 42.6 47.5
Orig 74.6 78.4 42.6 47.5
Role Role 74.5 78.2 43.5 47.5
Orig 74.5 78.2 43.5 47.0
Additionally, to investigate the impact of various prompt choices, we conduct new experiments with two alternative prompting techniques, following prior work planning and role-playing (Wang et al., 2023a; Kong et al., 2024). Specifically, we assess the robustness of ReFeri by varying prompts during both the generation stage and the verification stage by adapting different prompting styles (orig, plan, and role). For plan and role prompting at the generation, we newly sample five responses similar to Table 1. Table 4 shows the performance on MATH500 and GPQA under different configurations. A key observation is that verification performance remains highly consistent across different verification prompt styles for a given generation prompt style. For instance, on GPQA, performance for "plan $\rightarrow$ orig" and "plan $\rightarrow$ plan" conditions is identical (47.5 vs. 47.5), with similar consistency observed for the "role" condition (47.5 vs. 47.0). This indicates that ReFeri is inherently robust to variations in prompt formatting during the evaluation stage.
However, we also observe that the initial quality of the generated candidate set varies depending on the prompt style. For relatively simple tasks like MATH500, the quality of generated responses is similar across prompts. Conversely, on the more challenging GPQA, prompts offering structured guidance (e.g., plans or roles) tend to generate higher-quality seeds, reflected in slightly higher accuracy. Consequently, ReFeri performs better when the initial candidates are of higher quality.
@@
Estimation models and computational cost. To examine whether ReFeri is robust to the choice of estimation model $p_{\theta}$, we evaluate its performance using three LLMs with diverse scales and backbones: LLaMA-3.2-1B-Inst, Qwen-2.5-7B-Inst, and LLaMA-3.1-70B-Inst. The generation model is fixed (either GPT-4o-mini, GPT-4o, or LLaMA-3.1-8B), and we apply each estimation models to two tasks on MATH500 and GPQA. The average accuracy of three generation LLMs is presented in Figure 4 (Full results are in Appendix B.6). Here, ReFeri consistently improves Few-shot CoT across all settings, with an average gain of 4.9% on MATH500 and 5.1% on GPQA. Notably, the smallest model (LLaMA-3.2-1B) performs competitively, and even achieves competitive performance on MATH500. We attribute this to the relative simplicity of MATH benchmark, as recent small LLMs often exhibit reasonable performance; hence, they can make reliable likelihood estimates for selection. In contrast, GPQA requires more complex reasoning; therefore, using the large estimation model could be more beneficial. Indeed, LLaMA-3.1-70B achieves the best performance on this case. Despite these task-specific differences, the overall improvements are consistent across all estimation models. This suggests that the effectiveness of ReFeri primarily stems from its validation strategy with few-shot examples, rather than the specific choice of estimation model.
In addition, this consistent effectiveness offers better cost-accuracy trade-off. To show this, we conduct experiments using a small verifier for ReFeri (LLaMA-3.2-1B) and compared it against baselines that rely on a larger model (USC and CoT-WP with LLaMA-3.1-8B). Accuracy and latency per query (seconds per instance on a single GPU with identical configuration) are reported in Table 15. Here, ReFeri with a 1B estimator outperforms the strong 8B CoT-WP baseline while substantially reducing latency. For instance, on MATH500, ReFeri (1B) is approximately 60–65% faster than the CoT-WP (8B) baseline (e.g., 3.0s vs. 8.3s on MATH500). Furthermore, ReFeri exhibits robust performance regardless of estimation model size, whereas baselines often suffer significant degradation when scaled down. This demonstrates that combining ReFeri with a small-scale estimator provides a highly advantageous, delivering robust validation at remarkably low computational cost.
4 Related Works
Few-shot in-context learning of LLM. Few-shot in-context learning (ICL) revealed that LLMs can generalize to unseen tasks with just a handful of input-output demonstrations (Brown et al., 2020). To handle complex reasoning problems, chain-of-thought (CoT) prompting was proposed to append intermediate steps to the few-shot examples, leading to substantial gains in tasks such as arithmetic, commonsense reasoning, and symbolic manipulation (Wei et al., 2022; Fu et al., 2023; Jin et al., 2024). To further enhance ICL, various strategies have been developed to retrieve better examples using semantic similarity or entropy-based selection (Wu et al., 2023; Peng et al., 2024). However, some studies have shown that few-shot ICL does not always guarantee improvements. For instance, label shuffling or format changes can often leave performance unaffected (Min et al., 2022), and the performance gap between zero-shot and few-shot CoT is narrowing in several benchmarks as instruction tuning becomes more effective (Sprague et al., 2025). In particular, recent LLMs such as DeepSeek-R1, which are trained with reinforcement learning-based reasoning steps, sometimes even show performance degradation when few-shot CoT examples are added (Guo et al., 2025). Nonetheless, carefully selected demonstrations are still effective (Huang et al., 2024). For example, (Ge et al., 2025) show that few-shot examples can reduce overconfidence in multi-step reasoning, and (Yan et al., 2025) show that they help mitigate hallucinations and memory-based mistakes in complex tasks. These observations motivate us to go beyond using few-shot examples solely for generation, and recycling them to evaluate multiple LLM responses and to select the most promising one.
Selection of diverse LLM outputs. Due to the probabilistic nature of LLM decoding, LLM can provide diverse outputs for a single input, each reflecting different reasoning paths (Kadavath et al., 2022; Wang & Zhou, 2024; Qiu & Miikkulainen, 2024). To handle this variability, self-consistency (Wang et al., 2023b) samples $K$ independent reasoning paths and selects the majority answer to improve accuracy. However, it assumes that the model produces a single, well-formatted answer, and this assumption is often violated in open-ended tasks such as summarization or free-form dialogue (Stiennon et al., 2020; Salemi et al., 2024). Alternatively, recent Best-of-N approaches aim to directly select the best output among candidates, often using external verification models. For instance, Outcome Reward Models (ORMs) grade final outputs (Cobbe et al., 2021; Uesato et al., 2022), while Process Reward Models (PRMs) assess intermediate reasoning steps to provide finer supervision (Lightman et al., 2024; Wang et al., 2024a). Despite their successes, these models require large-scale, task-specific annotations or domain-specific checkers, limiting their scalability to new domains or unseen tasks. To eliminate the need for external verification models, prompting-based methods such as LLM-as-Judge ask LLM to evaluate its own outputs (Chen et al., 2023; Zheng et al., 2023). However, their effectiveness heavily depends on the model’s prior knowledge in the target domain. When this knowledge is lacking, these methods require additional fine-tuning with curated evaluation datasets for sufficient performance, which reintroduces the need for supervision (Yuan et al., 2024; Mahan et al., 2024; Zhang et al., 2025). In contrast, ReFeri is training-free and task-agnostic, offering a more scalable and generalizable approach by recycling a few-shot examples for verification.